Can you please help me with this question?
Two dice are rolled. What is the probability the sum of the two dice will be a prime number?

Found 2 solutions by stanbon, JeeJa:
Answer by stanbon(75887)

Two dice are rolled. What is the probability the sum of the two dice will be a prime number?
----
sum = 3: 1/2 , 2/1-----2 ways
sum = 5: 1/4,2/3,3/2,4/1----4 ways
sum = 7: 1/6,2/5,3/4,4/3,5/2,6/1----6 ways
---
Total # of ways to succeed: 2+4+6 = 12
---
# of possible sums: 36
---
P(prime sum) = 12/36 = 1/3
=============================

Answer by JeeJa(1)

prime nos are= 2, 3, 5, 7, 11
now,
sum=2 {(1,1)}
sum=3 {(1,2) (2,1)}
sum=5 {(1,4) (4,1) (2,3) (3,2)}
sum=7 {(1,6) (6,1) (2,5) (5,2) (3,4) (4,3)}
sum=11 {(5,6) (6,5)}
total=15
the no.of possible outcomes is 6*6=36
p(sum=prime no)= 15/36
therefore answer is 5/12